# Rethinking the Win Ratio: A Causal Framework for Hierarchical Outcome Analysis

### v1

@@ Abstract
Quantifying causal effects in the presence of complex and multivariate outcomes is a key challenge to evaluate treatment effects. For hierarchical multivarariates outcomes, the FDA recommends the Win Ratio and Generalized Pairwise Comparisons approaches Pocock et al. (2011); Buyse (2010). However, as far as we know, these empirical methods lack causal or statistical foundations to justify their broader use in recent studies. To address this gap, we establish causal foundations for hierarchical comparison methods. We define related causal effect measures, and highlight that depending on the methodology used to compute Win Ratios or Net Benefits of treatments, the causal estimand targeted can be different, as proved by our consistency results. Quite dramatically, it appears that the causal estimand related to the historical estimation approach can yield reversed and incorrect treatment recommendations in heterogeneous populations, as we illustrate through striking examples.
In order to compensate for this fallacy, we introduce a novel, individual-level yet identifiable causal effect measure that better approximates the ideal, non-identifiable individual-level estimand. We prove that computing Win Ratio or Net Benefits using a Nearest Neighbor pairing approach between treated and controlled patients, an approach that can be seen as an extreme form of stratification, leads to estimating this new causal estimand measure. We extend our methods to observational settings via propensity weighting, distributional regression to address the curse of dimensionality, and a doubly robust framework. We prove the consistency of our methods, and the double robustness of our augmented estimator. These methods are straightforward to implement, making them accessible to practitioners.
Key words.
Win Ratio, Hierarchical Outcomes, Multiple Outcomes, Randomized Control Trials, Observational data, Distributional Regression.
@@ 1.1 Contributions and outline of the paper
Finally, the last part of the paper is devoted to estimating the estimand we introduce. In Section 4, we extend the Nearest Neighbors approach to the observational setting by incorporating propensity weights. In order to face the fact that Nearest Neighbors may be slow to converge in the presence of high dimensional input features and to face their lack of robustness in the presence of missing data in the covariates, we introduce a distributional regression approach in Section 4.2 to estimate our new causal estimand. The distributional regression estimators we propose benefit from very simple implementations, and as such we believe they can be easily used by practitioners. We finally propose a doubly-robust estimator, that combines the strengths of the distributional regression approach and of the weighted Nearest Neighbors approach.
@@ 1.2 Related works
2 Causal inference framework for Win Ratio and generalized pairwise comparisons
@@ 2.2 Traditional Win Ratio, Net-Benefit and Mann-Whitney-Wilcoxon comparison estimators
    where $\sigma^{\star}(i):\mathcal{N}_{0}\to\mathcal{N}_{1}$ matches $i\in\mathcal{N}_{0}$ to the treated individual $j\in\mathcal{N}_{1}$ that has closest features in $\mathcal{X}$ i.e., $\sigma^{\star}(i)\in\operatorname{argmin}_{j\in\mathcal{N}_{1}}{{\left\|X_{i}-X_{j}\right\|}}^{2}$. $\mathcal{C}_{\mathrm{NN}}$ can also be generalized to $k-$Nearest Neighbors. For continuous features, Nearest Neighbors can be naturally applied after an eventual reweighting of the different features, to prevent from scale effects. For categorical features, Nearest Neighbors can be applied with one-hot encodings for instance, leading to Hamming-distances. In the presence of different categorical features of different importance, these can be weighted according to their importance.
@@ 4.1 An Inverse Propensity Weighting approach
    Finite and bounded second moment of propensity scores: lim $\limsup\mathbb{E}\left[\frac{1}{(1-\hat{\pi}(X_{i}))^{2}}\right]$ and lim $\limsup\mathbb{E}\left[\frac{1}{\hat{\pi}(X_{i})^{2}}\right]<\infty$.
@@ 4.2.1 The direct distributional regression estimator
$=\mathbb{E}\left[\mathbb{P}\left(T_{i}=0|X_{i}\right)\mathbb{E}\left[w(Y^{(X_{i})}(1)|Y_{i})|X_{i},Y_{i},T_{i}=0\right]+\mathbb{P}\left(T_{i}=1|X_{i}\right)\mathbb{E}\left[w(Y_{i}|Y^{(X_{i})}(0))|X_{i},Y_{i},T_{i}=1\right]\right]$
@@ 4.2.3 Distributional random forests and logistic regression for estimating $q_{1}(x,y)$
Our goal is to estimate $q_{1}(x,y)=\mathbb{P}\left(w(Y_{i}|y)|X_{i}=x,T_{i}=1\right)$ with distributional random forests (Cevid et al., 2022; Bénard et al., 2024). Let $\mathcal{H}$ a Hilbert space with a kernel $k(\cdot,\cdot)$ defined on $\mathcal{Y}\times\mathcal{Y}$ (usually the Gaussian kernel). For any probability measure $\mathbb{P}$ on $\mathcal{Y}$, let $\phi(\mathbb{P})=\mathbb{E}\left[k(Z,\cdot)\right]\in\mathcal{H}$, for $Z\sim\mathbb{P}$. Distributional random forests approximate $\mu(x,t)=\phi(\mathbb{P}_{Y|X=x,T=t})$ using splitting rules in the Hilbert space $\mathcal{H}$, and as such are simply generalization of vanilla random forests to Hilbert spaces. Then, Cevid et al. (2022) use the fact that for kernels such as the Gaussian kernel, learning kernel representations amounts to learning probability distributions. Thus, learning with samples $(X_{1},Y_{1}),\ldots,(X_{m},Y_{m})$, $\mu(x,t)$ is approximated via $\hat{\mu}_{m}(x,t)$ that takes the form:
@@
In practice, these steps are implemented in a very concise way444that will be released soon on github, using the following pseudo code to implement the distributional regression estimator (Equation 12).
@@ 5 Experiments on randomly generated data
We first start with experiments on random synthetic data. [^5] Further experiments on real world data — observational and trial data — will be added in a version that will be available shortly.
@@ Acknowledgements
M. Even acknowledges funding from PEPR Santé Numérique SMATCH [^6] https://pepr-santenum.fr/2023/11/07/smatch/. We would like to deeply thank Uri Shalit, Stefan Michiels and François Petit for very stimulating discussions in the writing of this papers, Dovid Parnas for suggesting Example 1 and for all the very the useful feedback, and Jeffrey Näf for the help in using the DRF package.

### v2

@@ Abstract
For hierarchical multivarariates outcomes, the FDA recommends the Win Ratio and Generalized Pairwise Comparisons approaches Pocock et al. (2011); Buyse (2010). However, as far as we know, these empirical methods lack causal or statistical foundations to justify their broader use in recent studies. To address this gap, we establish causal foundations for hierarchical comparison methods. We define related causal effect measures, and highlight that depending on the methodology used to compute Win Ratio, the causal estimand targeted can be different, as proved by our consistency results, which may then lead to reversed and incorrect treatment recommendations in heterogeneous populations, as we illustrate through striking examples.
In order to compensate for this fallacy, we introduce a novel, individual-level yet identifiable causal effect measure that better approximates the ideal, non-identifiable individual-level estimand. We prove that computing Win Ratio or Net Benefits using a Nearest Neighbor pairing approach between treated and controlled patients, an approach that can be seen as an extreme form of stratification, leads to estimating this new causal estimand measure. We extend our methods to observational settings via propensity weighting, distributional regression to address the curse of dimensionality, and a doubly robust framework. We prove the consistency of our methods, and the double robustness of our augmented estimator. These methods are straightforward to implement, making them accessible to practitioners. Finally, we validate our approach using synthetic data and on CRASH-3 (CRASH et al., 2019), a major clinical trial focused on assessing the effects of tranexamic acid in patients with traumatic brain injury.
Keywords: Hierarchical Outcomes Analysis, Multiple Outcomes, Randomized Control Trials, Observational data, Distributional Regression.
@@ 1.1 Contributions and outline of the paper
The last methodology part of the paper is devoted to estimating the estimand we introduce. In Section 4, we extend the Nearest Neighbors approach to the observational setting by incorporating propensity weights. In order to face the fact that Nearest Neighbors may be slow to converge in the presence of high dimensional input features and to face their lack of robustness in the presence of missing data in the covariates, we introduce a distributional regression approach in Section 4.2 to estimate our new causal estimand. The distributional regression estimators we propose benefit from very simple implementations, and as such we believe they can be easily used by practitioners. We propose a doubly-robust estimator, that combines the strengths of the distributional regression approach and of the weighted Nearest Neighbors approach.
We finally illustrate our methodologies on synthetically generated observational data and on a real world RCT. The CRASH-3 trial (Clinical Randomisation of an Antifibrinolytic in Significant Head Injury) was a large-scale, multi-center randomized controlled trial (RCT) that included over 12,000 patients in 175 hospitals across 29 countries. In particular, our Win Ratio study of the CRASH-3 dataset in Section 6 (CRASH et al., 2019) compares the existing Win Ratio methodologies with our proposed methods. We offer guidances to compute Win Ratio and confidence intervals for the estimators, and our illustration on the CRASH-3 dataset shows that while traditional approaches fail to offer statistically significant results (1 is in the computed confidence intervals), the methods we propose in this paper offer significant results (in favor of placebo treatment here), due to the fact that an individual-level estimand is being targeted, rather than a population-level one.
@@ 1.2 Related works
2 Causal inference framework
@@ 2.2 Traditional Win Ratio, Net-Benefit and Mann-Whitney-Wilcoxon comparison estimators
    where $\sigma^{\star}(i):\mathcal{N}_{0}\to\mathcal{N}_{1}$ matches $i\in\mathcal{N}_{0}$ to the treated individual $j\in\mathcal{N}_{1}$ that has closest features in $\mathcal{X}$ i.e., $\sigma^{\star}(i)\in\operatorname{argmin}_{j\in\mathcal{N}_{1}}{{\left\|X_{i}-X_{j}\right\|}}^{2}$. $\mathcal{C}_{\mathrm{NN}}$ can also be generalized to $k-$Nearest Neighbors.
Handling mixed factor and numerical features.
For continuous features, Nearest Neighbors can be naturally applied after an eventual reweighting of the different features, to prevent from scale effects. For categorical features, Nearest Neighbors can be applied with one-hot encodings for instance, leading to Hamming-distances. In the presence of different categorical features of different importance, these can be weighted according to their importance. In the CRASH-3 dataset that we study in Section 6, we handle mixed categorical and numerical features in two different ways. The first one is to consider the Mahalanobis distance, that naturally balances variabilities and scale of the different features. The second approach, that we believe to be even more robust and significant, uses a Factor Analysis of Mixed Data (FAMD). Since nearest neighbors algorithm relies on distance metrics (like Euclidean or Manahalahobis distances), it struggles with mixed data types. FAMD transforms both numerical and categorical variables into a common latent space, ensuring a more meaningful distance calculation. Furthermore, nearest neighbors algorithm suffers from high-dimensional data because distances become less meaningful in higher dimensions. FAMD captures the most important variations in fewer dimensions, improving NN’s effectiveness. This approached is detailed in Section 6 when studying the CRASH-3 data.
@@ 4.1 An Inverse Propensity Weighting approach
    Finite and bounded second moment of propensity scores:
    lim $\limsup\mathbb{E}\left[\frac{1}{(1-\hat{\pi}(X_{i}))^{2}}\right]\quad\text{and}\quad\limsup\mathbb{E}\left[\frac{1}{\hat{\pi}(X_{i})^{2}}\right]<\infty\,.$
@@ 4.2.1 The direct distributional regression estimator
$=\mathbb{E}\left[\mathbb{P}\left(T_{i}=0|X_{i}\right)\mathbb{E}\left[w(Y^{(X_{i})}(1)|Y_{i})|X_{i},Y_{i},T_{i}=0\right]\right]$
$\quad+\mathbb{E}\left[\mathbb{P}\left(T_{i}=1|X_{i}\right)\mathbb{E}\left[w(Y_{i}|Y^{(X_{i})}(0))|X_{i},Y_{i},T_{i}=1\right]\right]$
@@ 4.2.3 Distributional random forests and logistic regression for estimating $q_{1}(x,y)$
Our goal is to estimate
$q_{1}(x,y)=\mathbb{P}\left(w(Y_{i}|y)|X_{i}=x,T_{i}=1\right)\,,$
with distributional random forests (Cevid et al., 2022; Bénard et al., 2024). Let $\mathcal{H}$ a Hilbert space with a kernel $k(\cdot,\cdot)$ defined on $\mathcal{Y}\times\mathcal{Y}$ (usually the Gaussian kernel). For any probability measure $\mathbb{P}$ on $\mathcal{Y}$, let $\phi(\mathbb{P})=\mathbb{E}\left[k(Z,\cdot)\right]\in\mathcal{H}$, for $Z\sim\mathbb{P}$. Distributional random forests approximate $\mu(x,t)=\phi(\mathbb{P}_{Y|X=x,T=t})$ using splitting rules in the Hilbert space $\mathcal{H}$, and as such are simply generalization of vanilla random forests to Hilbert spaces. Then, Cevid et al. (2022) use the fact that for kernels such as the Gaussian kernel, learning kernel representations amounts to learning probability distributions. Thus, learning with samples $(X_{1},Y_{1}),\ldots,(X_{m},Y_{m})$, $\mu(x,t)$ is approximated via $\hat{\mu}_{m}(x,t)$ that takes the form:
@@
In practice, these steps are implemented in a very concise way, using the following pseudo code to implement the distributional regression estimator (Equation 12).
@@ 5 Experiments on randomly generated data
We first start with experiments on random synthetic data.
@@ 5.2 Misspecification and double robustness
6 Application to the CRASH-3 RCT
We finally illustrate our methodologies to perform a Win Ratio analysis of the CRASH-3 trial dataset, that comes from the CRASH-3 clinical trial CRASH et al. (2019), which studied the effects of tranexamic acid (TXA) in traumatic brain injury (TBI). In this section, we illustrate the different ways our methodologies can be used to derive Win Ratio estimates with confidence intervals. Overall, our conclusions are that different methods should be used for increased robustness of the results, since the properties of the estimands that are being targeted may differ.
6.1 Presentation of the dataset
Data description and preprocessing.
The CRASH-3 RCT contains information on 12, 737 patients. In order to have lighter computations, we chose to use only a random sample (without replacement) of 6, 000 patients for all our analysis. Missing data is imputed using mice (Van Buuren and Groothuis-Oudshoorn, 2011).
  1. 1.
    Patients covariates include: siteId (hospital identifier), sex (male/female), age (years), timeSinceInjury (hours since injury), sbpStatus (systolic blood pressure category), systolicBloodPressure (mmHg), gcsEyeOpening, gcsMotorResponse, gcsVerbalResponse (Glasgow Coma Scale scores), gcsTiming (time of GCS assessment), pupilReact (pupil reactivity), majorExtracranial (major extracranial injury), intraCranialBleeding (intracranial bleeding), epidural, subdural, subarachnoid, parenchymal, intraventricular (types of brain injuries), eligible (study eligibility), consent (study consent), eyeOpening, communicationAbility, motorResponse, feeding, toileting, grooming, levelOfFunctioning, employability, walking, washingDressing (functional outcomes), painDiscomfort, anxietyDepression, agitationAggression, fatigue (quality of life indicators), daysIcu (ICU stay duration), neuroHaemEvac (hematoma evacuation surgery), neuroOther (other neurosurgery), and estBloodLoss (estimated blood loss).
  2. 2.
    Patients are either assigned Placebo (6,321) or TXA (6,416 patients).
  3. 3.
    Primary outcomes are death events in the 28 days following trauma, that we encode as 1 or 0.
  4. 4.
    Secondary outcomes are vascular risks. We encode them as 1 (if there is at least one stroke, heart attack, pulmonary embolism or deep vein thrombosis) or 0 (if there are no such events).
  5. 5.
    Tertiary outcomes are the number of days the patient stayed in the hospital (censored at 28 days).
Computing the average treatment effect for each of these 3 outcomes lead respectively to the confidence intervals, where $Y_{1},Y_{2},Y_{3}$ are respectively our death, secondary effects and hospitalization duration outcomes:
$\mathbb{E}\left[Y_{1}(1)-Y_{1}(0)\right]\in[-0.0032,0.0033]\,,$
$\mathbb{E}\left[Y_{2}(1)-Y_{2}(0)\right]\in[-0.0021,0.0079]\,,$
and
$\mathbb{E}\left[Y_{3}(1)-Y_{3}(0)\right]\in[-0.2305,0.4512]\,,$
computed using the difference of means estimator. None of these intervals are significant. Here are some key conclusions from the study made by CRASH et al. (2019), to further contextualize the dataset:
  1. 1.
    TXA reduces the risk of death due to TBI, but only if given within 3 hours of injury. The earlier TXA is given, the greater the benefit. No significant benefit if given after 3 hours.
  2. 2.
    TXA did not increase the risk of Stroke, Heart attack, Pulmonary embolism, Deep vein thrombosis (DVT).
  3. 3.
    Greatest Benefit for Mild-to-Moderate TBI. TXA had the most impact on patients with mild-to-moderate TBI (Glasgow Coma Scale 9-15). No significant survival benefit in severe TBI (GCS $\leqslant$ 8), possibly because of the high fatality rate.
  4. 4.
    No Increase in Disability. TXA did not increase the number of survivors with severe disability. Patients who survived had similar functional outcomes to those in the placebo group.
  5. 5.
    Safe and Cost-Effective. TXA is cheap and widely available, making it a practical treatment for emergency trauma care. Safe for use in pre-hospital settings and emergency departments.
The final takeaway is that TXA is an effective, safe, and low-cost intervention that can save lives in TBI when given early (within 3 hours). However, it does not help much in severe TBI and must be administered as soon as possible after injury.
6.2 The different methodologies used
We compared the following methodologies in Figure 7:
  1. 1.
    Traditional win ratio, computed using all pairs of the dataset using the WINS package (Cui and Huang, 2021).
  2. 2.
    Stratified Win Ratio (Dong et al., 2018): we stratify according to the time since injury the patient received the treatment (TXA or placebo). We computed the median of time since injury, and made 2 stratas: patients below and above this median.
  3. 3.
[…]
  4. 4.
    Another version of nearest neighbors: Optimal Matchings, that solve (in the case where $|\mathcal{N}_{1}|\leqslant|\mathcal{N}_{0}|$
    $\min_{\sigma:\mathcal{N}_{1}\to\mathcal{N}_{0}\text{ injective}}\sum_{i\in\mathcal{N}_{1}}{{\left\|X_{\sigma(i)}-X_{i}\right\|}}^{2}\,.$
    The difference with Nearest Neighbors as studied in our paper is that two treated units cannot be assigned to the same control unit. We use this on both the full covariates and on the features in the FAMD latent space.
  5. 5.
    Distributional Random Forests, as described in Section 4.2. We estimate the Win Proportion (obtained with $w(y,y^{\prime})=\mathds{1}_{{\left\{y\succ y^{\prime}\right\}}}$) and the Loss Proportion (obtained with $w(y,y^{\prime})=\mathds{1}_{{\left\{y\preceq y^{\prime}\right\}}}$), and estimate the Win Ratio as the ratio between win et loss proportions.
  6. 6.
    Our doubly robust approach, as described in Section 4.2. For the DRF and AIPW DRF, forests are taken with 1000 trees.
For confidence intervals, we either use those given by the WINS package for computing win ratios (that use asymptotic gaussian approximations), or use bootstrapping with 1,000 bootstraps for distributional regression approaches. Gaussian approximations for Win Ratio consist of outputting confidence intervals that are of the form:
Win $\text{Win Ratio}\in\frac{\#\text{Wins}\pm 1,96\sqrt{\#\text{Wins}}}{\#\text{Losses}\pm 1,96\sqrt{\#\text{Losses}}}\,.$
Figure 7: Win ratio computed on the CRASH-3 dataset using different methodologies.
6.3 Analysis of Figure 7.
The four pairing methods (complete pairings, stratified pairings according to the time since injury, nearest neighbors and optimal pairing) are compared in Figure 7.
Win Ratio with complete and stratified pairings.
These two pairings do not show significant results: 1 is in the confidence intervals, meaning that the Win Ratio analysis with these methods cannot conclude for or against treatment.
Nearest neighbors and optimal pairings on the covariates with the Mahanalobis distance.
Nearest neighbors have the same behavior if performed on the covariates. However, optimal pairings offer significant results against arm 1 in favor of placebo.
Nearest neighbors and optimal pairings on the FAMD latent space.
However, when performed on the FAMD latent space, Win Ratio computed using nearest neighbors and optimal pairings are both significant (in favor of placebo). This suggests that computing matchings and distances in different ways may lead to slightly different results. We believe that in our case, due to mixed categorical and numerical features, the FAMD approach with distances in this latent space seems to be the right approach.
Distributional regression approaches.
The boxplots of the two distributional regression approaches (DRF and AIPW DRF) are however much larger, reminiscent of the larger variability of forest based approaches. As such, results for the DRF approach are not significant. However, the augmented approach, AIPW DRF, offers significant results (in favor of placebo) despite its large variance, illustrating its double robustness properties.
Conclusion from this study.
Our Win Ratio analysis of the CRASH-3 study suggestsa preference over placebo rather than over treatment. This is not in contradiction with the conclusion of the CRASH-3 study, that were in favor of treatment only on a subpopulation. ATEs computed on the whole population are indeed not statistically significant, as showed above. This suggests is the strength of our methodologies, that can lead to significant discoveries, in scenarios where traditional or stratified Win Ratio fail to do so. Indeed, as shown in the simple synthetic example in Examples 1 and 2, in the presence of heterogeneity one must be very careful at which estimand is being targeted. Further investigations would be necessary to draw clinical conclusions from a Win Ratio analysis of the CRASH-3 dataset.
@@ Acknowledgements
M. Even acknowledges funding from PEPR Santé Numérique SMATCH [^4] https://pepr-santenum.fr/2023/11/07/smatch/. We would like to deeply thank Uri Shalit, Stefan Michiels and François Petit for very stimulating discussions in the writing of this paper, Dovid Parnas for suggesting Example 1 and for all the very the useful feedback, and Jeffrey Näf for the help in using the DRF package.